# Zo3T: Zero-shot 3D-Aware Trajectory-Guided image-to-video generation via Test-Time Training

### v2

@@ 4 Experiments
Evaluation Metrics. We conduct a comprehensive evaluation using both quantitative metrics and a user study. For quantitative assessment, we measure Video Quality via FID, FVD, and four key VBench (Huang et al. 2024) metrics (subject consistency, background consistency, aesthetic quality, and imaging quality), and Motion Control Accuracy using Object Motion Conformity (ObjMC) (Namekata et al. 2024)—the Euclidean distance between target and actual trajectories. In parallel, our user study involved six expert evaluators who assessed 50 videos generated from a custom dataset of images, each with a predefined trajectory.
Evaluation Datasets. For a fair comparison, we adopt the open-source dataset used in (Wu et al. 2024). Recognizing that many baseline I2V models excel on common, in-distribution trajectories (Wang et al. 2024a), we augment this set by applying a “mirroring” transformation to most paths, resulting in a more challenging set of 143 samples featuring more diverse and less common motion paths. Following the protocols in (Wu et al. 2024), FID and FVD metrics are computed on the VIPSeg (Miao et al. 2022) dataset.
@@
We perform a comprehensive evaluation by comparing Zo3T with SOTA supervised and zero-shot baselines. For supervised baselines, we select DragNUWA (Yin et al. 2023), DragAnything (Wu et al. 2024) and LeviTor (Wang et al. 2025b). The zero-shot methods include SG-I2V (Namekata et al. 2024), ObjCtrl-2.5D (Wang et al. 2024a), and an adapted version of FreeTraj (Qiu et al. 2024) for our I2V task, following the procedure in (Namekata et al. 2024). For a fair evaluation, all video outputs are resized to a 320×576 resolution to align with the supervised models.
@@
To validate the contribution of each core component, we conduct both quantitative (Table 2) and qualitative (Figure 6) ablation studies.
@@ 7 Experiment Setup Details
7:  $(\mathbf{z}_{t}^{\text{opt}},\Delta\theta_{\text{LoRA}})\leftarrow$ Iteratively optimize $(\mathbf{z}_{t},\Delta\theta_{\text{LoRA}})$ by minimizing $\mathcal{J}_{\text{TTT}}$ (Eq. 3).
8:  $\mathbf{z}_{t}^{*}\leftarrow\mathcal{M}_{fuse}(\mathbf{z}_{t}^{\text{opt}},\mathbf{z}_{t})$ using Eq. 6.
@@
13:  $\mathbf{d}_{t}^{*}\leftarrow$ Iteratively optimize $\mathbf{d}$ from $\mathbf{d}_{0}$ by minimizing $\mathcal{J}_{\text{guide}}$ (Eq. 4).
@@
For FVD calculation, we follow the protocol established in prior works (Wu et al. 2024; Namekata et al. 2024). All videos, including our generated results and the ground-truth videos from the VIPSeg dataset (Miao et al. 2022), are resized to $\times256$. We employ the ViCLIP (Wang et al. 2023) model as the feature extractor. The FVD is computed as the Fréchet distance between the distributions of real and generated video features:
@@
Following the approach in DragNUWA (Yin et al. 2023), we evaluate motion fidelity using the Object Motion Consistency (ObjMC) metric. This metric is defined as the average per-frame Euclidean distance between the tracked trajectory and the ground-truth trajectory. The trajectory in the generated video is estimated using Co-Tracker (Karaev et al. 2024).

### v3

@@ 4 Experiments
Evaluation Metrics. We conduct a comprehensive evaluation using both quantitative metrics and a user study. For quantitative assessment, we measure Video Quality via FID, FVD, and four key VBench (huang2023vbench) metrics (subject consistency, background consistency, aesthetic quality, and imaging quality), and Motion Control Accuracy using Object Motion Conformity (ObjMC) (namekata2024sg)—the Euclidean distance between target and actual trajectories. In parallel, our user study involved six expert evaluators who assessed 50 videos generated from a custom dataset of images, each with a predefined trajectory.
Evaluation Datasets. For a fair comparison, we adopt the open-source dataset used in (wu2024draganything). Recognizing that many baseline I2V models excel on common, in-distribution trajectories (objctrl2.5d), we augment this set by applying a “mirroring” transformation to most paths, resulting in a more challenging set of 143 samples featuring more diverse and less common motion paths. Following the protocols in (wu2024draganything), FID and FVD metrics are computed on the VIPSeg (vipseg) dataset.
@@
We perform a comprehensive evaluation by comparing Zo3T with SOTA supervised and zero-shot baselines. For supervised baselines, we select DragNUWA (yin2023dragnuwa), DragAnything (wu2024draganything) and LeviTor (levitor). The zero-shot methods include SG-I2V (namekata2024sg), ObjCtrl-2.5D (objctrl2.5d), and an adapted version of FreeTraj (qiu2024freetraj) for our I2V task, following the procedure in (namekata2024sg). For a fair evaluation, all video outputs are resized to a 320×576 resolution to align with the supervised models.
@@
We validate each core component via quantitative (Table 2) and qualitative (Figure 6) ablations.
@@ 5 Conclusion
Acknowledgements
This work was partly supported by Shenzhen Key Laboratory of next generation interactive media innovative technology (No:ZDSYS20210623092001004).
@@ 7 Experiment Setup Details
7:  $(\mathbf{z}_{t}^{\text{opt}},\Delta\theta_{\text{LoRA}})\leftarrow$ Iteratively optimize $(\mathbf{z}_{t},\Delta\theta_{\text{LoRA}})$ by minimizing $\mathcal{J}_{\text{TTT}}$.
8:  $\mathbf{z}_{t}^{*}\leftarrow\mathcal{M}_{fuse}(\mathbf{z}_{t}^{\text{opt}},\mathbf{z}_{t})$.
@@
13:  $\mathbf{d}_{t}^{*}\leftarrow$ Iteratively optimize $\mathbf{d}$ from $\mathbf{d}_{0}$ by minimizing $\mathcal{J}_{\text{guide}}$.
@@
For FVD calculation, we follow the protocol established in prior works (wu2024draganything; namekata2024sg). All videos, including our generated results and the ground-truth videos from the VIPSeg dataset (vipseg), are resized to $\times256$. We employ the ViCLIP (internvid) model as the feature extractor. The FVD is computed as the Fréchet distance between the distributions of real and generated video features:
@@
Following the approach in DragNUWA (yin2023dragnuwa), we evaluate motion fidelity using the Object Motion Consistency (ObjMC) metric. This metric is defined as the average per-frame Euclidean distance between the tracked trajectory and the ground-truth trajectory. The trajectory in the generated video is estimated using Co-Tracker (cotracker).